OpenStudy (anonymous):

What is the solution set for |3x – 24| 27? x 1 or x 17 x –1 and x 17 x 17 or x –17 x –1 and x 17

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x - 24 = \pm 27 }\)

OpenStudy (anonymous):

This isn't helping... : (

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x - 24 = 27 }\) or \( \color{Black}{\Rightarrow 3x - 24 = -27 }\)

OpenStudy (anonymous):

So its c

Parth (parthkohli):

Nope nope

OpenStudy (anonymous):

B?

Parth (parthkohli):

Yep :)